OpenStudy (anonymous):

use the given zero to find all the zeros of the function. f(x)=x^3+9x^2+27x+35 given x=-2+√3

OpenStudy (anonymous):

\[x^3+9 x^2+27 x+35=(x+5) \left(x^2+4 x+7\right) \]\[x^2+4 x+7=0\]\[x=-2\pm i \sqrt{3} \]